OpenStudy (jaylelile):

Please help! Richard has investment totaling $4,000 in two accounts. One is a savings account paying 3% annual interest and the other a bond paying 5% annual interest. If the annual interest from the two investments was $180, how much did he have invested at each rate? b. Write the system of equations to represent this situation c. solve the system

OpenStudy (kropot72):

Let the amount invested at 3% be A, and let the amount invested at 5% be B. Then A + B = 4000 .....................(1) When we consider the interest we can write: A * 0.03 + B * 0.05 = 180 .......(2) Now you need to solve the system of equations (1) and (2) to find the values of A and B.

OpenStudy (kropot72):

@jaylelile Are you there?

OpenStudy (jaylelile):

yes just trying to understand...

OpenStudy (kropot72):

Which part are you having trouble understanding?

OpenStudy (jaylelile):

how to solve the system

OpenStudy (kropot72):

If we look at the first equation A + B = 4000 .....................(1) it can be rearranged to get B in terms of A, as follows: B = 4000 - A .................(3) If you now replace B in equation (2) by (4000 - A) you can solve the resulting equation to find the value of A.